OpenStudy (anonymous):

The line that passes through (-2, 4) and is parallel to x - 2y = 6

OpenStudy (anonymous):

pareral lines have same gradient so x - 2y = 6 2y = x - 6 y = (1/2)x - 3 so the gradient is 1/2 so the parallel line has the same gradientso Y = (1/2)X + c then plug -2,4 (as the line passes through that point) to that equation to get c 4 = (1/2) * (-2) + c c = 5 so the equation is Y = (1/2)x + 5 2y = x + 10 x - 2y - 10 = 0